OpenStudy (ilovecake):

What is the value of the variable? I will give medal 13c = 39 c =

OpenStudy (anonymous):

c=3 13 times 3= 39

OpenStudy (midhun.madhu1987):

Divide both sides by 13.

OpenStudy (ilovecake):

I still don't get it

OpenStudy (triciaal):

you have 13 c and you want 1 c so you need to divide by 13 since you have an equation whatever you do to one side you do to the other to keep the balance so you divide 39 by 13 as well

OpenStudy (triciaal):

13 *c = 39 13*c/13 = 39/13 c =

OpenStudy (ilovecake):

Do I divide 13 by 39???

OpenStudy (ilovecake):

c = 3?

OpenStudy (ilovecake):

@triciaal

OpenStudy (ilovecake):

ugh

OpenStudy (ilovecake):

someone

OpenStudy (ilovecake):

help

OpenStudy (anonymous):

c=3

OpenStudy (anonymous):

yah welcs

OpenStudy (ilovecake):

ok thx

OpenStudy (anonymous):

np

OpenStudy (triciaal):

yes c = 3 ; I was in the middle of another problem